# Gaussian Processes for Time Series with Lead-Lag Effects with application to biology data

### v1

@@ Abstract
Investigating the relationship, particularly the lead-lag effect, between time series is a common question across various disciplines, especially when uncovering biological process. However, analyzing time series presents several challenges. Firstly, due to technical reasons, the time points at which observations are made are not at uniform inintervals. Secondly, some lead-lag effects are transient, necessitating time-lag estimation based on a limited number of time points. Thirdly, external factors also impact these time series, requiring a similarity metric to assess the lead-lag relationship. To counter these issues, we introduce a model grounded in the Gaussian process, affording the flexibility to estimate lead-lag effects for irregular time series. In addition, our method outputs dissimilarity scores, thereby broadening its applications to include tasks such as ranking or clustering multiple pair-wise time series when considering their strength of lead-lag effects with external factors. Crucially, we offer a series of theoretical proofs to substantiate the validity of our proposed kernels and the identifiability of kernel parameters. Our model demonstrates advances in various simulations and real-world applications, particularly in the study of dynamic chromatin interactions, compared to other leading methods.
@@ 1 Introduction
The lead-lag effect, a widespread phenomenon where changes in one time series (the leading series) influence another time series (the influenced follower series) after a certain delay. This phenomenon is prevalent in a variety of fields, such as climate science (De Luca and Pizzolante, 2021), healthcare (Zhu and Gallego, 2021; Runge et al., 2019), and economics (Wang et al., 2017; Skoura, 2019). For example, Harzallah and Sadourny (1997) explores the lead-lag relationship between the Indian summer monsoon and various climate variables, such as sea surface temperature, snow cover, and geopotential height. However, this dynamic, rooted in causal regulatory relationships, is notably observed in biological processes, which serves as the primary inspiration for the proposed method. The lead-lag effect manifests itself in various processes, including development (Gerstein et al., 2010; Ding and Bar-Joseph, 2020; Strober et al., 2019), disease-associated genetic variation (Krijger and De Laat, 2016), and responses to various interventions and treatments (Faryabi et al., 2008; Lu et al., 2021).
@@
Addressing this challenge, our study focuses on the multi-omics time series data presented by Reed et al. (2022). This dataset includes measurements of enhancer activity, and promoter transcription, and 3D chromatin structure at seven time points, from 0 to 6 hours (a later time point is excluded from this analysis). These data reveal complex dynamic patterns, as depicted in Figure 1. Our goal is to identify functional human enhancer-promoter pairs from over 20,000 candidate pairs, estimate the time lag, and quantify the extent of the lead-lag effect between enhancer activity and gene expression over time.
@@
Beyond quantifying the lead-lag effect between pairwise time series, there is often a desire to estimate time lag considering the irregularities in time series data, one approach is to first impute values at regular time points, then apply methods like TLCC to calculate the time lag. Common techniques for this separate modeling of each time series include using splines (Rehfeld et al., 2011) or separate Gaussian processes (SGP, Li et al. (2021)). However, such methods do not leverage shared information across time series, potentially leading to sub-optimal results. Hierarchical GP models such as Magma (Leroy et al., 2022, 2023) integrate shared information across multiple time series but are primarily designed for prediction and clustering individual time series. This is in contrast to our primary objective, which is to cluster pairwise time series by a measure of similarity. More direct approaches for irregular time series include the use of the negative and positive lead-lag estimator (NAPLES, Ito and Nakagawa (2020)) and Lead-Lag method (Hoffmann et al., 2013), which utilize the Hayashi–Yoshida covariance estimator to estimate a fixed time-lag. Alternatively, multi-task GPs through latent variable models introduce additional variables to account for temporal discrepancies, such as GPLVM (Dürichen et al., 2014), which may limit flexibility due to the presumed independence between tasks and temporal sequences (also known as separable kernels). Moreover, more complex misalignment techniques (Kazlauskaite et al., 2019; Mikheeva et al., 2022) might not perform well with limited time observations for biology data where simpler linear mismatch assumptions suffice, and they can face operational challenges like Cholesky decomposition failures. Table 1 compares the aforementioned methods, highlighting their suitability for limited or irregular time points, their ability to quantify the magnitude of lead-lag effects, and estimate time lags:
@@
SGP $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
Splines $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
GPLVM $*{}^{*}$ $*{}^{*}$
MAGMA $*{}^{*}$ $*{}^{*}$
@@
Note: $*{}^{*}$ means yes if followed by TLCC or DTW.
@@
All proofs are in the Appendix, along with additional experimental details. All codes are available at https://github.com/Wancen/GPlag/.
2 Models
@@ 2.2 Pairwise Time Series Gaussian Process Model
The proof is given in Appendix B. Since the key idea in the above theorem and corollary is motivated by Gneiting (2002), we call the kernel family constructed from Corollary 1 the Gneiting family. In particular, Gneiting enumerates four candidates for $\phi$ and three for $\psi$ in Gneiting (2002). Exploring each amalgamation of $\phi$ and $\psi$ families allows for further tailoring by adjusting parameters intrinsic to the family, culminating in the desired kernel. Furthermore, more kernels can be constructed by combining kernels from Gneiting family and other simple kernels. Analogues of the RBF, exponential kernel, and Matérn kernel in Euclidean space are provided in the following definitions.
@@
As an illustration, LRBF can be derived by setting $\phi(t)=e^{-bt}$ and $\psi(t)=a^{2}t+1$ in Corollary 1; LExp is the product of a kernel in Gneiting family with $\phi(t)=e^{-bt^{1/2}}$, $\psi(t)\equiv 1$, and a simple kernel $\frac{1}{\mathrm{\bf 1}_{\{l\neq l^{\prime}\}}a^{2}+1}$; LMat is the product of a kernel in Gneiting family with $\phi(t)=\frac{(bt^{1/2})^{\nu}}{2^{\nu-1}\Gamma(\nu)}K_{\nu}(bt^{1/2})$, $\psi\equiv 1$, and a simple kernel $\frac{1}{(\mathrm{\bf 1}_{\{l\neq l^{\prime}\}}a^{2}+1)^{\nu+1/2}}$. Additional candidate kernels are presented in the Appendix A, along with a detailed explanation on how they are constructed.
@@ 2.3 Interpretation of parameters
The interpretability of kernel parameters is one of the major advantages of GP models. In the kernels defined above, $\sigma^{2}$, the spatial variance, controls the point-wise variance; $b$, the lengthscale parameter, governs the temporal dependency; the smoothness parameter $\nu$ determines the smoothness of the sample path. In this paper, we assume $\nu$ to be given, which is a common assumption in GP literature (Rasmussen, 2004). More importantly, $a$, is called the dissimilarity parameter, where a smaller $a$ signifies a lesser degree of dissimilarity, which corresponds to a stronger lead-lag effect. For two distinct time series $l\neq l^{\prime}$, when $a=\infty$, the covariance $K((t,l),(t^{\prime},l^{\prime}))=0$, indicating independence between time series, or no lead-lag effect at all. On the other hand, when $a=0$, $K((t,1),(t^{\prime},2))=\sigma^{2}e^{-b(t-t^{\prime}-s)^{2}}$, implying that the two time series are the same up to a time lag, the strongest lead-lag effect. The interpretation of $a$ in GPlag parallels measures lead-lag effect in other classical methods like DTW and TLCC: $a=0$ in GPlag aligns with a loss of 0 in DTW and a correlation of 1 in TLCC to signify the strongest lead-lag effect. Similarly, $a=\infty$ in GPlag corresponds to a loss of infinity in DTW and a correlation of 0 in TLCC for the weakest effect. The time lag parameter $s$ measures the time lag between two time series. In the next subsection, we discuss the implementation of GPlag and the inference of these parameters, especially $a$ and $s$. Theoretical support for the identifiability of $a$ and $s$ is provided in Section 3.
@@ 2.4 Implementation
where $Y=[y_{1},\cdots,y_{n}]^{\top}$ denotes the vector of observations, and $\Sigma$ is the $(n_{1}+n_{2})\times(n_{1}+n_{2})$ covariance matrix specified in Equations 2, 3 and 4. Optimization methods such as L-BFGS-B (Byrd et al., 1995) and Adam  (Kingma and Ba, 2014) are widely used. The L-BFGS-B algorithm allows for setting lower and upper bounds on the parameters, enhancing the stability of the numeric optimization. The parameters $a,\leavevmode\nobreak\ b,\leavevmode\nobreak\ \sigma^{2},\leavevmode\nobreak\ \tau^{2}$ are constrained to be positive and $s$ can be set to a reasonable range depending on specific cases. For the initial values of $b,\leavevmode\nobreak\ \sigma^{2},\leavevmode\nobreak\ \tau^{2}$, we suggest running a standard GP regression first and using the estimates as our initial values. For $s$, we recommend taking the estimate from TLCC, and we recommend $a=1$ as the initialization. The illustration of MLE process is provided in Algorithm 1.
The extension to Bayesian inference is straightforward: we can simply sample from the posterior $\pi(a,b,s,\sigma^{2},\tau^{2}|\{t_{i},c_{i},y_{i}\}_{i=1}^{n})\propto\pi(a,b,s,\sigma^{2},\tau^{2})p(Y|\{y_{i},c_{i},t_{i}\}_{i=1}^{n},a,b,s,\sigma^{2},\tau^{2})$, where $\pi(a,b,s,\sigma^{2},\tau^{2})$ is the joint prior distribution. We suggest using independent inverse Gamma priors for $a,\leavevmode\nobreak\ b,\leavevmode\nobreak\ \sigma^{2},\leavevmode\nobreak\ \tau^{2}$ and a Gaussian prior for $s$. The algorithm is depicted in Algorithm 3 and has been implemented in Python using GPyTorch (Gardner et al., 2018) and the Pyro (Bingham et al., 2019) package. This implementation enables rapid, fully Bayesian inference and takes advantage of GPU acceleration.
@@ 2.5 Extension to multi-time series
The positive definite, symmetric, and triangle inequality constraints on $a_{ll^{\prime}}$ ensures the positive definiteness of $K$, while the constraint on $s_{1}$ is due to our treating the first time series as the baseline, meaning that the time lag is relative to the first time series. The inference is similar to the previous case, with the only difference being that the optimization method involved in MLE needs to consider these additional constraints. In other words, it is a constrained optimization problem. Nevertheless, it is not intractable, as all the constraints on $a_{ll^{\prime}}$ are linear and can be accommodated by existing optimization methods, including L-BFGS-B and Adam, as illustrated in Algorithm 2 with three time series.
@@ 3 Identifiability of GPlag Kernel Parameter
In this section, we prove that the parameters $A$ and $S$ in GPlag model are identifiable, allowing them to be used to as measures of the lead lag effect between time series and the time lag. We begin by introducing the necessary definitions.
@@
As a result, it suffices to show the following necessary and sufficient conditions for two GPlags being equivalent:
@@
The proof is given in Appendix C. Although $\sigma^{2}$ and $b$ in GPlag are not identifiable, the proposed interpretation of the model does not rely on these two parameters. Instead, the parameter $A$ is interpreted as a measure of dissimilarity between time series under a time lag represented by the parameter $S$.
@@ 4 Simulation studies
The main focus of our experiments was to illustrate that GPlag could accurately estimate both the time lag and the extent of lead-lag effects. This capability aided in the ranking or clustering of multiple time series that existed in a lead-lag relationship with a target time series. We demonstrated the theoretical properties of GPlag using synthetic data, as these properties were best illustrated under a fully controlled environment. All code and additional implementation details were available in the supplement.
@@
Our results showed that the MLE of $a$ and $s$ were very close to their truth among all three kernels when $n=100$. This was consistent with the theoretical analysis in Theorem 3 (Figure 2 A, B) that GPlag could accurately recover the true lead-lag effects. Compared to Lead-Lag, GPlag exhibited similar or smaller variance, specifically when the kernel was Matérn with $\nu=3/2$. Additionally, we performed simulations with a varying number of time points and found that the variance of the MLEs of $a$ and $s$ decreased as the sample size increased, which empirically confirmed the consistency of the MLEs (Figure 2 C).
Prediction. We assessed the accuracy of GPlag’s predictions by generating data from two different processes. In the first, we used the kernel LExp in Equation 3 with parameters $b=1,\leavevmode\nobreak\ a=0.3,\leavevmode\nobreak\ s=2$, and $\sigma^{2}=4$. In the second, we generated pairwise time series from linear functions with non-Gaussian noise (t-distributed): $y=2t+3+5\epsilon_{1}$ and $y=2(t-20)+3+5\epsilon_{2}$ where $t=0,1,2,\ldots,100$. The noise term, $\epsilon_{1},\epsilon_{2}$, were drawn from a t-distribution with 5 degrees of freedom. We then randomly selected $50\%$ of the data as training data for model fitting and used the remaining $50\%$ as testing data for prediction in both settings. We applied the best linear unbiased prediction (BLUP, see the supplement for more details) and used the mean squared error (MSE) on the testing set as the metric to evaluate the prediction accuracy.
Among all methods, GPlag achieved the highest log-likelihood and smallest MSE (Figure 3 A-B, Figure 6). The advantage of GPlag over SGP and splines was that it modeled two time series together which allowed for sharing of information between time series through $a$. When $a\rightarrow\infty$, it was equivalent to SGP. While a small estimated $a$ indicated that two time series were similar, and the covariance function then allowed for more information to be shared between them. Although Magma adopted a hierarchical GP to specifically model the mean function for irregular data, it did not consider time lag information across time series.
@@
Nine pairs of time series were fitted by GPlag with the LExp kernel. We observed that three pairs of time series in each column with the same $k$ were grouped together (Figure 3 D), and $a$ increased as $k$ increased. This implied that GPlag could accurately rank the degree of lead-lag effects among multiple pairs of time series based on the parameter $a$, through precise estimation of the time lag parameter $s$. A larger estimate of $a$ indicated a stronger dissimilarity between pairs of time series. To further quantify the clustering performance of lead-lag effects among the nine pairs of time series, we used the Adjusted Rand Index (ARI, Hubert and Arabie (1985)) and Normalized Mutual Information (NMI, Strehl and Ghosh (2002)) as evaluation metrics. To define a cluster from the pairwise dissimilarity parameter (correlation coefficients for TLCC and distance loss for DTWs), we used K-means (Shannon, 1948) with 3 clusters. As the focus of this work was on determining pairwise dissimilarity rather than the clustering algorithm, we opted for a simple clustering method. GPlag achieved the highest ARI and NMI in clustering the lead-lag effects among 9 pairwise time series (Table 2, column 2-3).
@@
Figure 4: Fitting performance on three time series generated from LExp with b = 1, $a_{12}=a_{13}=a_{23}=1$, $s_{2}=2,s_{3}=4$. (A) MLE of $a_{12},a_{13},a_{23},s_{1},s_{2}$. The true parameter values are represented by the dashed horizontal lines. (B) Prediction MSE on the held-out dataset, from 100 replicates.
5 Real-world application: Dynamic chromatin interactions
In this section, we applied GPlag to a time-resolved multi-omics dataset involving enhancers and genes in human cells activated by interferon gamma from a resting state (Reed et al., 2022). Originally, the dataset comprised over 20,000 pairs. Each pair consisted of one time series for each enhancer and likewise one time series for each gene, at seven irregularly-spaced time points (0, 0.5, 1, 1.5, 2, 4, and 6 hours after a perturbation). The time series captured gene expression measurements for genes and histone tail acetylation measurements (H3K27ac), a measure of activity for the enhancers, both approximately variance stabilized. It is widely believed that the alterations in acetylation occur prior to the changes in expression of the regulated genes (Reed et al., 2022). However, focusing only on pairs that became functional after stimulation, we applied a filter. This filter, based on a minimum-maximum difference of stabilized value across the timescale greater than 1, yielded 4,776 dynamic pairs of time series. Subsequently, the objective of our analysis, after estimating the model parameters of GPlag for these 4,776 pairs, was to identify which subset represented functional enhancer-gene pairs, i.e. the enhancer regulates the gene though with an unknown lag before expression changes occur for the gene.

### v2

@@ Abstract
Investigating the relationship, particularly the lead-lag effect, between time series is a common question across various disciplines, especially when uncovering biological process. However, analyzing time series presents several challenges. Firstly, due to technical reasons, the time points at which observations are made are not at uniform intervals. Secondly, some lead-lag effects are transient, necessitating time-lag estimation based on a limited number of time points. Thirdly, external factors also impact these time series, requiring a similarity metric to assess the lead-lag relationship. To counter these issues, we introduce a model grounded in the Gaussian process, affording the flexibility to estimate lead-lag effects for irregular time series. In addition, our method outputs dissimilarity scores, thereby broadening its applications to include tasks such as ranking or clustering multiple pair-wise time series when considering their strength of lead-lag effects with external factors. Crucially, we offer a series of theoretical proofs to substantiate the validity of our proposed kernels and the identifiability of kernel parameters. Our model demonstrates advances in various simulations and real-world applications, particularly in the study of dynamic chromatin interactions, compared to other leading methods.
@@ 1 Introduction
The lead-lag effect, a widespread phenomenon where changes in one time series (the leading series) influence another time series (the influenced follower series) after a certain delay. This phenomenon is prevalent in a variety of fields, such as climate science (De Luca and Pizzolante, 2021), healthcare (Runge et al., 2019; Zhu and Gallego, 2021; Feng et al., 2021), economics (Wang et al., 2017; Skoura, 2019), and finance (Hoffmann et al., 2013; Bacry et al., 2013; Da Fonseca and Zaatour, 2017; Ito and Sakemoto, 2020). For example, Harzallah and Sadourny (1997) explores the lead-lag relationship between the Indian summer monsoon and various climate variables, such as sea surface temperature, snow cover, and geopotential height. However, this dynamic, rooted in causal regulatory relationships, is notably observed in biological processes, which serves as the primary inspiration for the proposed method. The lead-lag effect manifests itself in various processes, including development (Gerstein et al., 2010; Ding and Bar-Joseph, 2020; Strober et al., 2019), disease-associated genetic variation (Krijger and De Laat, 2016), and responses to various interventions and treatments (Faryabi et al., 2008; Lu et al., 2021).
@@
Addressing this challenge, our study focuses on the multi-omics time series data presented by Reed et al. (2022). This dataset includes measurements of enhancer activity, and promoter transcription, and 3D chromatin structure at seven time points, from 0 to 6 hours (a later time point is excluded from this analysis). These data reveal complex dynamic patterns, as depicted in Figure 1. Our goal is to identify functional human enhancer-promoter pairs from over 20,000 candidate pairs, estimate the time lag, and quantify the extent of the lead-lag effect between enhancer activity and gene expression over time. The biological motivation for this goal is to uncover a relationship in genomics time series for a subset of enhancer-gene pairs, in particular those in which the enhancer state changes after cell activation and causes a subsequent change in expression of a target gene. Measuring the lag between enhancer activity and gene expression over time for these candidate enhancer-promoter pairs helps in understanding the sequence of regulatory events set off by activation, and how the time scale of gene regulation may vary across the genome.
@@
Beyond quantifying the lead-lag effect between pairwise time series, there is often a desire to estimate time lag considering the irregularities in time series data, one approach is to first impute values at regular time points, then apply methods like TLCC to calculate the time lag. Common techniques for this separate modeling of each time series include using splines (Rehfeld et al., 2011) or separate Gaussian processes (SGP, Li et al. (2021)). However, such methods do not leverage shared information across time series, potentially leading to sub-optimal results. Hierarchical GP models such as Magma (Leroy et al., 2022, 2023) integrate shared information across multiple time series but are primarily designed for prediction and clustering individual time series. This is in contrast to our primary objective, which is to cluster pairwise time series by a measure of similarity. More direct approaches for irregular time series include the use of the negative and positive lead-lag estimator (NAPLES, Ito and Nakagawa (2020)) and Lead-Lag method (Hoffmann et al., 2013), which utilize the Hayashi–Yoshida covariance estimator to estimate a fixed time-lag. Alternatively, multi-task GPs through latent variable models introduce additional variables to account for temporal discrepancies such as GPLVM (Dürichen et al., 2014), which may limit flexibility due to the presumed independence between tasks and temporal sequences (also known as separable kernels). Moreover, more complex misalignment techniques (Kazlauskaite et al., 2019; Mikheeva et al., 2022) might not perform well with limited time observations for biology data where simpler linear mismatch assumptions suffice, and they can face operational challenges like Cholesky decomposition failures. Table 1 compares the aforementioned methods, highlighting their suitability for limited or irregular time points, their ability to quantify the magnitude of lead-lag effects, and estimate time lags:
@@
SGP X X X X
Splines X X X X
GPLVM X X
MAGMA X X
@@
Note: X means yes if followed by TLCC or DTW.
@@
2 Methods
@@ 2.2 Pairwise Time Series Gaussian Process Model
The proof is given in Section A.1. Since the key idea in the above theorem and corollary is motivated by Gneiting (2002), we call the kernel family constructed from Corollary 1 the Gneiting family. In particular, Gneiting enumerates four candidates for $\phi$ and three for $\psi$ in Gneiting (2002). Exploring each amalgamation of $\phi$ and $\psi$ families allows for further tailoring by adjusting parameters intrinsic to the family, culminating in the desired kernel. Furthermore, more kernels can be constructed by combining kernels from Gneiting family and other simple kernels. Analogues of the RBF, exponential kernel, and Matérn kernel in Euclidean space are provided in the following definitions.
@@
As an illustration, LRBF can be derived by setting $\phi(t)=e^{-bt}$ and $\psi(t)=a^{2}t+1$ in Corollary 1; LExp is the product of a kernel in Gneiting family with $\phi(t)=e^{-bt^{1/2}}$, $\psi(t)\equiv 1$, and a simple kernel $\frac{1}{\mathrm{\bf 1}_{\{l\neq l^{\prime}\}}a^{2}+1}$; LMat is the product of a kernel in Gneiting family with $\phi(t)=\frac{(bt^{1/2})^{\nu}}{2^{\nu-1}\Gamma(\nu)}K_{\nu}(bt^{1/2})$, $\psi\equiv 1$, and a simple kernel $\frac{1}{(\mathrm{\bf 1}_{\{l\neq l^{\prime}\}}a^{2}+1)^{\nu+1/2}}$. Common choices for $\nu$ are $1/2,3/2,5/2$, where the Bessel function $K_{\nu}$ admits a closed form, similar to the standard Matérn kernel, with the specific choice depending on the data. Additional candidate kernels are presented in the Appendix A Appendix A, along with a detailed explanation on how they are constructed.
@@ 2.3 Interpretation of parameters
The interpretability of kernel parameters is one of the major advantages of GP models. In the kernels defined above, $\sigma^{2}$, the spatial variance, controls the point-wise variance; $b$, the lengthscale parameter, governs the temporal dependency; the smoothness parameter $\nu$ determines the smoothness of the sample path. In this paper, we assume $\nu$ to be given, which is a common assumption in GP literature (Rasmussen, 2004). More importantly, $a$, is called the dissimilarity parameter, where a smaller $a$ signifies a lesser degree of dissimilarity, which corresponds to a stronger lead-lag effect. For two distinct time series $l\neq l^{\prime}$, when $a=\infty$, the covariance $K((t,l),(t^{\prime},l^{\prime}))=0$, indicating independence between time series, or no lead-lag effect at all. On the other hand, when $a=0$, $K((t,1),(t^{\prime},2))=\sigma^{2}e^{-b(t-t^{\prime}-s)^{2}}$, implying that the two time series are the same up to a time lag, the strongest lead-lag effect. The interpretation of $a$ in GPlag parallels measures lead-lag effect in other classical methods like DTW and TLCC: $a=0$ in GPlag aligns with a loss of 0 in DTW and a correlation of 1 in TLCC to signify the strongest lead-lag effect. Similarly, $a=\infty$ in GPlag corresponds to a loss of infinity in DTW and a correlation of 0 in TLCC for the weakest effect. The time lag parameter $s$ measures the time lag between two time series. In the next subsection, we discuss the implementation of GPlag and the inference of these parameters, especially $a$ and $s$. Further visualization of sample paths of above kernels with different combinations of parameters are provided in Figure S5-Figure S7. Theoretical support for the identifiability of $a$ and $s$ is provided in Section 3.
@@ 2.4 Implementation
where $Y=[y_{1},\cdots,y_{n}]^{\top}$ denotes the vector of observations, and $\Sigma$ is the $(n_{1}+n_{2})\times(n_{1}+n_{2})$ covariance matrix specified in Equations 2, 3 and 4. Optimization methods such as L-BFGS-B (Byrd et al., 1995) and Adam  (Kingma and Ba, 2014) are widely used. The L-BFGS-B algorithm allows for setting lower and upper bounds on the parameters, enhancing the stability of the numeric optimization. The parameters $a,\leavevmode\nobreak\ b,\leavevmode\nobreak\ \sigma^{2},\leavevmode\nobreak\ \tau^{2}$ are constrained to be positive and $s$ can be set to a reasonable range depending on specific cases. For the initial values of $b,\leavevmode\nobreak\ \sigma^{2},\leavevmode\nobreak\ \tau^{2}$, we suggest running a standard GP regression first and using the estimates as our initial values. For $s$, we recommend taking the estimate from TLCC, and we recommend $a=1$ as the initialization. The illustration of MLE process is provided in Appendix A Algorithm 1.
The extension to Bayesian inference is straightforward: we can simply sample from the posterior $\pi(a,b,s,\sigma^{2},\tau^{2}|\{t_{i},c_{i},y_{i}\}_{i=1}^{n})\propto\pi(a,b,s,\sigma^{2},\tau^{2})p(Y|\{y_{i},c_{i},t_{i}\}_{i=1}^{n},a,b,s,\sigma^{2},\tau^{2})$, where $\pi(a,b,s,\sigma^{2},\tau^{2})$ is the joint prior distribution. We suggest using independent inverse Gamma priors for $a,\leavevmode\nobreak\ b,\leavevmode\nobreak\ \sigma^{2},\leavevmode\nobreak\ \tau^{2}$ and a Gaussian prior for $s$. The algorithm is depicted in Algorithm 3 in Section A.3 and has been implemented in Python using GPyTorch (Gardner et al., 2018) and the Pyro (Bingham et al., 2019) package. This implementation enables rapid, fully Bayesian inference and takes advantage of GPU acceleration.
@@ 2.5 Extension to multi-time series
The positive definite, symmetric, and triangle inequality constraints on $a_{ll^{\prime}}$ ensures the positive definiteness of $K$, while the constraint on $s_{1}$ is due to our treating the first time series as the baseline, meaning that the time lag is relative to the first time series. The inference is similar to the previous case, with the only difference being that the optimization method involved in MLE needs to consider these additional constraints. In other words, it is a constrained optimization problem. Nevertheless, it is not intractable, as all the constraints on $a_{ll^{\prime}}$ are linear and can be accommodated by existing optimization methods, including L-BFGS-B and Adam, as illustrated in Appendix A Algorithm 2 with three time series.
@@ 3 Identifiability of GPlag Kernel Parameter
In this section, we prove that the parameters $A$ and $S$ in GPlag model are identifiable, allowing them to be used to as measures of the lead lag effect between time series (measured by $A$) and the time lag (measured by $S$). Here, we assume all parameters are positive, finite, real numbers, and the domain is fixed (also known as infill domain). The fixed domain assumption is reasonable for multi-omics time series, as the dynamics of gene regulation in such experiments occur during a fixed window of time. Investigators will decide, based on available budget, the time points needed to capture the dynamics during, say activation, or differentiation. If allocated more budget, they will then fill in the gaps between existing observations, to have better resolution over the same time period. We begin by introducing the necessary definitions.
@@
As a result, it suffices to show the following necessary and sufficient conditions for two GPlags being equivalent, which suggests identifiable parameters for further interpretation:
@@
The proof is given in Appendix A.2. Although $\sigma^{2}$ and $b$ in GPlag are not identifiable, the proposed interpretation of the model does not rely on these two parameters. Instead, the parameter $A$ is interpreted as a measure of dissimilarity between time series under a time lag represented by the parameter $S$.
@@ 4 Simulation studies
The main focus of our experiments was to illustrate that GPlag could accurately estimate both the time lag and the extent of lead-lag effects. This capability aided in the ranking or clustering of multiple time series that existed in a lead-lag relationship with a target time series. We demonstrated the theoretical properties of GPlag using synthetic data, as these properties were best illustrated under a fully controlled environment. All code and additional implementation details were available in the Section B.1.
@@
Our results showed that the MLE of $a$ and $s$ were very close to their truth among all three kernels when $n=100$. This was consistent with the theoretical analysis in Theorem 3 (Figure 2 A, B) that GPlag could accurately recover the true lead-lag effects. Compared to Lead-Lag, GPlag exhibited similar or smaller variance, specifically when the kernel was Matérn with $\nu=3/2$. Additionally, we performed simulations with a varying number of time points and found that the variance of the MLEs of $a$ and $s$ decreased as the sample size increased, which empirically confirmed the consistency of the MLEs (Figure 2 C). Plots for other parameters of secondary interest, including $\sigma^{2}$, $b$, and $\sigma^{2}b^{2\nu}$, are provided in Figure S1. We also performed simulations with different settings of $b=0.3,5,10$, and found that the consistency of the MLEs for $a$ and $s$ was maintained across these different values of $b$ (See Figure S2 for more details).
Prediction. We assessed the accuracy of GPlag’s predictions by generating data from two different processes. In the first, we used the kernel LExp in Equation 3 with parameters $b=1,\leavevmode\nobreak\ a=0.3,\leavevmode\nobreak\ s=2$, and $\sigma^{2}=4$. In the second, we generated pairwise time series from linear functions with non-Gaussian noise (t-distributed): $y=2t+3+5\epsilon_{1}$ and $y=2(t-20)+3+5\epsilon_{2}$ where $t=0,1,2,\ldots,100$. The noise term, $\epsilon_{1},\epsilon_{2}$, were drawn from a t-distribution with 5 degrees of freedom. We then randomly selected $50\%$ of the data as training data for model fitting and used the remaining $50\%$ as testing data for prediction in both settings. We applied the best linear unbiased prediction (BLUP, see Section B.1 for more details) and used the mean squared error (MSE) on the testing set as the metric to evaluate the prediction accuracy.
Among all methods, GPlag achieved the highest log-likelihood and smallest MSE (Figure 3 A-B, Figure S3). The advantage of GPlag over SGP and splines was that it modeled two time series together which allowed for sharing of information between time series through $a$. When $a\rightarrow\infty$, it was equivalent to SGP. While a small estimated $a$ indicated that two time series were similar, and the covariance function then allowed for more information to be shared between them. Although Magma adopted a hierarchical GP to specifically model the mean function for irregular data, it did not consider time lag information across time series.
@@
Nine pairs of time series were fitted by GPlag with the LExp kernel. We observed that three pairs of time series in each column with the same $k$ were grouped together (Figure 3 D), and $a$ increased as $k$ increased. This implied that GPlag could accurately rank the degree of lead-lag effects among multiple pairs of time series based on the parameter $a$, through precise estimation of the time lag parameter $s$. A larger estimate of $a$ indicated a stronger dissimilarity between pairs of time series. To further quantify the clustering performance of lead-lag effects among the nine pairs of time series, we used the Adjusted Rand Index (ARI, Hubert and Arabie (1985)) and Normalized Mutual Information (NMI, Strehl and Ghosh (2002)) as evaluation metrics. To define a cluster from the pairwise dissimilarity parameter (correlation coefficients for TLCC and distance loss for DTWs), we used K-means (Shannon, 1948) with 3 clusters. As the focus of this work was on determining pairwise dissimilarity rather than the clustering algorithm, we opted for a simple clustering method. GPlag achieved the highest ARI and NMI in clustering the lead-lag effects among 9 pairwise time series (Table 1, column 2-3). We also tested GPlag using the LRBF and LMat kernels, and found that ARI and NMI scores remained at 1, consistent with the LExp results. The inferred $a$ values are shown in Figure S4, further indicating that the results are not sensitive to the choice of kernel.
@@
Figure 4: Fitting performance on three time series generated from LExp with b = 0.3, $a_{12}=a_{13}=a_{23}=1$, $s_{2}=2,s_{3}=4$. (A) MLE of $a_{12},a_{13},a_{23},s_{1},s_{2}$. The true parameter values are represented by the dashed horizontal lines. (B) Prediction MSE on the held-out dataset, from 100 replicates.
5 Genomics application: Dynamic chromatin interactions
In this section, we applied GPlag to a time-resolved multi-omics dataset involving enhancers and genes in human cells activated by interferon gamma from a resting state (Reed et al., 2022). Originally, the dataset comprised over 20,000 pairs. Each pair consisted of one time series for each enhancer and likewise one time series for each gene, at seven irregularly-spaced time points (0, 0.5, 1, 1.5, 2, 4, and 6 hours after a perturbation). The time series captured gene expression measurements for genes and histone tail acetylation measurements (H3K27ac), a measure of activity for the enhancers, both approximately variance stabilized. It is widely believed that the alterations in acetylation occur prior to the changes in expression of the regulated genes (Reed et al., 2022). We focused only on pairs that became functional after stimulation, by applying a filter on the total range of stabilized values across the timescale, yielding 4,776 dynamic pairs of enhancer-gene time series. Subsequently, the objective of our analysis, after estimating the model parameters of GPlag for these 4,776 pairs, was to identify the subset of all possible enhancer-gene pairs that represent functional enhancer-gene pairs, i.e. the enhancer responds to cell activation and regulates the expression of the gene, after some time lag. Many types of time profiles can be captured by such a time series experiment, e.g. increasing across the range, decreasing, or increasing then decreasing, etc. Particular functional enhancer-gene pairs will have subtle differences in the dynamics of the coordinated change in activity and expression, compared to other functional pairs, which we hope to capture with our GPlag model.